Circular No. 7998

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 T5 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery obser-

vation given below), posted on the NEO Confirmation Page due to unusual motion, has been found to be cometary in appearance. D. T. Durig and J. A. K. Blackwood, University of the South, Sewanee, TN, report that CCD images taken with a 0.30-m reflector on Oct. 18.4 UT show the object to be diffuse with a hint of a tail 12′′–15′′ long in p.a. 330^{o} and a soft coma of diameter 5′′–8′′. Images taken by G. Hug, Eskridge, KS, also with a 0.30-m reflector at about the same time, show a nearly stellar coma with extension ∼ 10′′–15′′ long in p.a. ∼ 280^{o}.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 5.38947  4^{h}44^{m}17.^{s}33  −12^{o}15′26′′.7  18.4

Astrometry (Oct. 5–22), the following elliptical orbital elements, and an ephemeris appear on MPEC 2002-U20.

T = 2003 July 1.7593 TT  ω = 327.^{o}4640
e = 0.438402  Ω = 123.2783 }2000.0
q = 3.924792 AU  i = 30.9153
a = 6.988610 AU n^{o} = 0.0533480 P = 18.475 years

SATELLITES OF JUPITER The IAU Working Group on Planetary System Nomenclature has ap-

proved the following new designations and names:

Jupiter XVII Callirrhoe = S/1999 J 1
Jupiter XVIII Themisto = S/1975 J 1 = S/2000 J 1
Jupiter XIX  Magaclite = S/2000 J 8
Jupiter XX Taygete = S/2000 J 9
Jupiter XXI  Chaldene = S/2000 J 10
Jupiter XXII Harpalyke = S/2000 J 5
Jupiter XXIII Kalyke  = S/2000 J 2
Jupiter XXIV Iocaste  = S/2000 J 3
Jupiter XXV Erinome = S/2000 J 4
Jupiter XXVI Isonoe  = S/2000 J 6
Jupiter XXVII Praxidike = S/2000 J 7
2002 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green